Discourses in Mathematics and Its Applications, No. 4 1995

ISBN 0-9630728-3-8

Heat Kernel Techniques and Quantum Gravity

Edited by S. A. Fulling

Preface

The Conference on Heat Kernel Techniques and Quantum Gravity was held in Winnipeg, Manitoba, Canada, on the campus of the University of Manitoba during August 2–6, 1994. The conference received financial assistance from the National Science and Engineering Research Council of Canada, the Winnipeg Institute for Theoretical Physics, and the International Science Foundation. Additional financial support and facilities were provided by the Faculty of Science and the Department of Physics at the University of Manitoba.

The Organizing Committee for the conference consisted of

A. O. Barvinsky  (University of Alberta)
S. A. Fulling  (Texas A&M University)
G. Kunstatter  (University of Winnipeg)
T. A. Osborn  (University of Manitoba)
G. A. Vilkovisky  (Lebedev Physical Institute)

They are grateful to the staff members of the Department of Mathematics of Texas A&M University who helped to produce and distribute this volume: Brent Burton, Karola Feltz, Jan Want, and especially technical typist Robin Campbell.